# Progressive Compression with Universally Quantized Diffusion Models

### v1

@@ Abstract
Diffusion probabilistic models have achieved mainstream success in many generative modeling tasks, from image generation to inverse problem solving. A distinct feature of these models is that they correspond to deep hierarchical latent variable models optimizing a variational evidence lower bound (ELBO) on the data likelihood. Drawing on a basic connection between likelihood modeling and compression, we explore the potential of diffusion models for progressive coding, resulting in a sequence of bits that can be incrementally transmitted and decoded with progressively improving reconstruction quality. Unlike prior work based on Gaussian diffusion or conditional diffusion models, we propose a new form of diffusion model with uniform noise in the forward process, whose negative ELBO corresponds to the end-to-end compression cost using universal quantization. We obtain promising first results on image compression, achieving competitive rate-distortion-realism results on a wide range of bit-rates with a single model, bringing neural codecs a step closer to practical deployment.
@@ 1 Introduction
Ho et al. (2020); Theis et al. (2022) first proposed a progressive compression codec based on a diffusion model and demonstrated its strong potential for data compression. Such a progressive codec is desirable as it allows us to decode data reconstructions from partial bit-streams, starting from lossy reconstructions at low bit-rates to perfect (lossless) reconstructions at high bit-rates, all with a single model. The decoder’s ability to decode intermediate reconstructions without having to wait for all bits to be received is a highly useful feature available in many traditional codecs, such as JPEG. The use of diffusion models has the additional advantage that we can, in theory, obtain perfectly realistic reconstructions (Theis et al., 2022), even at ultra-low bit-rates. Unfortunately, the proposed method requires the communication of Gaussian samples across many steps, which remains intractable because of an exponential runtime complexity (Goc & Flamich, 2024).
In this work, we take first steps towards a diffusion-based progressive codec that is computationally tractable. The key idea is to replace Gaussian distributions in the forward process with suitable uniform distributions and accordingly adjust the reverse process distributions. These modifications allow the application of universal quantization (Zamir & Feder, 1992) for simulating the uniform noise channel, avoiding the intractability of Gaussian channel simulation in (Theis et al., 2022).
Specifically, our contributions are as follows
@@
    We investigate design choices of UQDM, specifying its forward and reverse processes largely by matching the moments of those in Gaussian diffusion, and obtain the best results when learning the reverse-process variance as inspired by (Nichol & Dhariwal, 2021).
@@ 2 Background
Relative Entropy Coding (REC) deals with the problem of efficiently communicating a single sample from a target distribution $q$ using a coding distribution $p$. Suppose two parties in communication have access to a common “prior” distribution $p$ and pseudo-random number generators with a common seed; a Relative Entropy Coding (REC) method (Flamich et al., 2020; Theis & Ahmed, 2022) allows the sender to transmit a sample ${\mathbf{z}}\sim q$ using close to $\mathrm{KL}(q\,\|\,p)$ bits, up to a logarithmic overhead. At a high level, a typical REC method works as follows. The sender generates a (possibly large) number of candidate ${\mathbf{z}}$ samples from the prior $p$,
@@
A major challenge of REC algorithms is that their computational complexity generally scales exponentially with the amount of information being communicated (Agustsson & Theis, 2020; Goc & Flamich, 2024). As an example, the MRC algorithm (Cuff, 2008; Havasi et al., 2018) draws $M$ candidate samples and selects $K\in\{1,2,,...,M\}$ with a probability proportional to the importance weights, $\nicefrac{{q({\mathbf{z}}_{n})}}{{p({\mathbf{z}}_{n})}},n=1,...,M$; similarly to importance sampling, $M$ needs to be on the order of $2^{\operatorname{KL}(q\|p)}$ for ${\mathbf{z}}_{K}$ to be (approximately) a fair sample from $q$, thus requiring a number of drawn samples that scales exponentially with the relative entropy $\operatorname{KL}(q\|p)$ (and a coding cost of $\log M\approx\operatorname{KL}(q\|p)$ bits). The exponential complexity prevents, e.g., naively communicating the entire latent tensor ${\mathbf{z}}$ in a Gaussian VAE for lossy compression, as the relative entropy $\mathrm{KL}(q({\mathbf{z}}|{\mathbf{x}})\,\|\,p({\mathbf{z}}))$ easily exceeds thousands of bits, even for a small image. This difficulty can be partly remedied by performing REC on sub-problems with lower dimensions (Flamich et al., 2020; 2022) for which computationally viable REC algorithms exist (Flamich et al., 2024; Flamich, 2024), but at the expense of worse bitrate efficiency due to the accumulation of codelength overhead across the dimensions.
@@
In the above, the optimal entropy coding distribution $\mathbb{P}(K|U^{\prime}=u^{\prime})$ is obtained by discretizing $p_{\tilde{Y}}=p_{Y}\star\mathcal{U}(-\nicefrac{{\Delta}}{{2}},\nicefrac{{\Delta}}{{2}})$ on a grid of width $\Delta$ and offset by $U^{\prime}=u^{\prime}$ (Zamir & Feder, 1992), where $\star$ denotes convolution. If the true $p_{\tilde{Y}}$ is unknown, we can replace it with a surrogate density model $f_{\theta}(\tilde{y})$ during entropy coding and incur a higher coding cost,
@@ 5 Experiments
When considering the progressive compression performance of VDM and UQDM, we consider three kinds of progressive reconstructions (from ${\mathbf{z}}_{t}$): denoise, where $\hat{{\mathbf{x}}}=\hat{{\mathbf{x}}}_{\theta}({\mathbf{z}}_{t};t)$ is the prediction from the denoising network; ancestral, where $\hat{{\mathbf{x}}}\sim p({\mathbf{x}}|{\mathbf{z}}_{t})$ is drawn by ancestral sampling; and flow-based where $\hat{{\mathbf{x}}}\sim p({\mathbf{x}}|{\mathbf{z}}_{t})$ is computed deterministically using the probability flow ODE in (Theis et al., 2022). In Gaussian diffusion, the probability flow ODE produces the same trajectory of marginal distributions as ancestral sampling. In the case of UQDM, we apply the same update equations and observe similar benefits, likely due to the continuous-time equivalence of the underlying processes of UQDM and VDM. See Section B.3 for details. Note that DiffC-A and DiffC-F (Theis et al., 2022) directly correspond to our VDM results with ancestral and flow-based reconstructions.
In all experiments involving VDM and UQDM, we always use the same U-net architecture for both, except UQDM uses twice as many output dimensions for both the denoising prediction and learned reverse-process variance (see Sec. 3). We refer to Appendix Sec. C for further experiment details.
@@ 5.1 Swirl Toy Data
Finally, we examine the quality of unconditional samples from UQDM with varying $T$. Although our earlier results indicate worse compression performance for $T>5$, Figure 7 shows that UQDM’s sample quality monotonically improves with increasing $T$. We conjecture this is because the reverse process of UQDM may also converge to that of VDM in the continuous-time limit.
@@ 5.2 CIFAR10
Next, we apply our method to natural images. We start with the CIFAR10 dataset containing $32\times 32$ images. We train a baseline VDM model with a smaller architecture than that used by Kingma et al. (2021), converging to around 3 bits per dimension. We use the noise schedule $\sigma_{t}^{2}=\mathrm{\sigma}(\gamma_{t})$ where $\gamma_{t}$ is a linear in $t$ with learned endpoints $\gamma_{T}$ and $\gamma_{0}$. For our UQDM model we empirically note that $T=4$ yields the best trade-off between bit-rates and reconstruction quality. We train our model end-to-end on the progressive coding NELBO eq. (1) with learned reverse-process variances.
@@
As shown in Figure 3, we consistently outperform both JPEG and JPEG2000 over all bit-rates and metrics. Even though BPG, a competitive non-progressive codec optimized for rate-distortion performance, achieves better reconstruction fidelity (as measured in PSNR) in the low bit-rate regime, our method closely matches BPG in realism (as measured in FID) and even beats BPG in PSNR at higher bit-rates. The theoretical performance of compression with Gaussian diffusion (e.g., VDM) (Theis et al., 2022), especially with a high number of steps such as $T=1000$, is computationally infeasible, both due to the large number of neural function evaluations required, and due the intractable runtime of REC algorithms in the Gaussian case. Still, for reference we report theoretical results both for $T=1000$ and $T=20$, where the latter uses a smaller and more practical number of diffusion/progressive reconstruction steps.
@@ 6 Discussion
Future work may further improve our approach to close the performance gap to Gaussian diffusion (Theis et al., 2022); the latter represents the ideal lossy compression performance under a perfect realism constraint on, e.g., approximately Gaussian-distributed data. To this end, potential improvements include better reconstruction methods than ancestral sampling, such as the ODE-based method considered in (Theis et al., 2022), or exploring different parameterizations of the forward and reverse process with better theoretical properties. Finally, we expect further improvement in computation efficiency and scalability when combining our method with ideas such as latent diffusion (Rombach et al., 2022), distillation (Sauer et al., 2024), or consistency models (Song et al., 2023).
@@ Acknowledgments
Justus Will and Yibo Yang acknowledge support from the HPI Research Center in Machine Learning and Data Science at UC Irvine. Stephan Mandt acknowledges support from the National Science Foundation (NSF) under an NSF CAREER Award IIS-2047418 and IIS-2007719, the NSF LEAP Center, by the Department of Energy under grant DE-SC0022331, the IARPA WRIVA program, the Hasso Plattner Research Center at UCI, the Chan Zuckerberg Initiative, and by gifts from Qualcomm and Disney. We thank Kushagra Pandey for feedback on the manuscript.
@@ A.2 Uniform (Ours)
${\mathbf{z}}_{t}=b_{T|t}\,{\mathbf{z}}_{T}+c_{T|t}\,{\mathbf{x}}+\underbrace{\sum_{v=t+1}^{T}\sqrt{12}\delta_{v|t}\mathbf{u}_{v}}_{\coloneqq{\boldsymbol{\omega}}_{t}},$
@@
$\delta_{v|t}:=\beta_{v|v-1}\prod_{j=t+1}^{v-1}b_{j|j-1}=\frac{\sigma_{t}^{2}}{\alpha_{t}}\sqrt{\mathrm{SNR}(v-1)-\mathrm{SNR}(v)}.$
@@
$\mathrm{Var}\left(\boldsymbol{\omega}_{t}\right)$ $=\sum_{v=t+1}^{T}\delta_{v|t}^{2}=\frac{\sigma_{t}^{4}}{\alpha^{2}_{t}}[\mathrm{SNR}(t)-\mathrm{SNR}(T)]=\beta^{2}_{T|t},$
@@ A.3 Convergence to the Gaussian case
We show that both parameterizations are equivalent in the continuous-time limit. To allow comparison across different number of steps $T$, we suppose that $\alpha_{t}$ and $\sigma_{t}$ are obtained from continuous-time schedules $\alpha(\cdot):[0,1]\to\mathbb{R}^{+}$ and $\sigma(\cdot):[0,1]\to\mathbb{R}^{+}$ (which were fixed ahead of time), such that $\alpha_{t}:=\alpha(t/T)$ and $\sigma_{t}:=\sigma(t/T)$ for $t=0,\dots,T$, for any choice of $T$. We further assume that the continuous-time signal-to-noise ratio $\mathrm{snr}(\cdot)=\alpha(\cdot)^{2}/\sigma(\cdot)^{2}$ is strictly monotonically decreasing.
We start by showing that our $q({\mathbf{z}}_{t}|{\mathbf{z}}_{T},{\mathbf{x}})$ converges to the corresponding Gaussian distribution in VDM in the continuous-time limit, which in turn implies the convergence of our $q({\mathbf{z}}_{t}|{\mathbf{x}})$ to the corresponding Gaussian distribution in VDM, for each fixed $t$.
@@
$\frac{1}{s}\sum_{i=1}^{n}\operatorname{\mathbb{E}}\left[X_{ni}^{2}\mathbf{1}\{\left|X_{ni}\right|\geq{\epsilon}\}\right]\xrightarrow{n\to\infty}0$
holds for all ${\epsilon}>0$. In this case, $s=\mathrm{Var}\left(Z_{n}\right)=\sigma_{0}^{4}/\alpha_{0}^{2}(\mathrm{snr}(0)-\mathrm{snr}(1))=\beta^{2}_{T|0}$. The condition holds trivially as by construction $P(\left|X_{ni}\right|\geq\sqrt{3}\,\sigma_{ni})=0$ and for every ${\epsilon}>0$ there exists $N_{\epsilon}$ with ${\epsilon}>\sqrt{3}\,\sigma_{ni}$ for all $i$ and $n>N_{\epsilon}$ as $\max_{i\in\{1,\dots n\}}\sigma_{ni}^{2}\to 0$. The statement follows for $t=0$ as $Z_{n}\sim\boldsymbol{\omega}_{t}|_{T=n}$, and analogously for arbitrary $t$ by considering $\sigma_{ni}^{2}\coloneqq 12\sigma_{t}^{4}/\alpha_{t}^{2}(\mathrm{snr}(t+(i-1)(1-t)/n)-\mathrm{snr}(t+i(1-t)/n))$. ∎
@@
As $q({\mathbf{z}}_{T}|x)=\mathcal{N}(\alpha_{T}{\mathbf{x}},\sigma^{2}_{T}\mathbf{I})=\mathcal{N}(\mathbf{0},\mathbf{I})$ does not depend on $T$, the joint distribution $q({\mathbf{z}}_{t},{\mathbf{z}}_{T}|x)=q({\mathbf{z}}_{t}|{\mathbf{z}}_{T},{\mathbf{x}})q({\mathbf{z}}_{T}|x)$ converges in distribution, which in turn implies convergence of $q({\mathbf{z}}_{t}|x)$. The statement then follows from the identity

### v2

@@ Abstract
Diffusion probabilistic models have achieved mainstream success in many generative modeling tasks, from image generation to inverse problem solving. A distinct feature of these models is that they correspond to deep hierarchical latent variable models optimizing a variational evidence lower bound (ELBO) on the data likelihood. Drawing on a basic connection between likelihood modeling and compression, we explore the potential of diffusion models for progressive coding, resulting in a sequence of bits that can be incrementally transmitted and decoded with progressively improving reconstruction quality. Unlike prior work based on Gaussian diffusion or conditional diffusion models, we propose a new form of diffusion model with uniform noise in the forward process, whose negative ELBO corresponds to the end-to-end compression cost using universal quantization. We obtain promising first results on image compression, achieving competitive rate-distortion and rate-realism results on a wide range of bit-rates with a single model, bringing neural codecs a step closer to practical deployment.
000
Preprint. Under Review.
@@ 1 Introduction
Ho et al. (2020); Theis et al. (2022) first suggested a progressive compression method based on an unconditional diffusion model and demonstrated its strong potential for data compression. Such a progressive codec is desirable as it allows us to decode data reconstructions from partial bit-streams, starting from lossy reconstructions at low bit-rates to perfect (lossless) reconstructions at high bit-rates, all with a single model. The ability to decode intermediate reconstructions without having to wait for all bits to be received is a highly useful feature present in many traditional codecs, such as JPEG. The use of diffusion models has the additional advantage that we can, in theory, obtain perfectly realistic reconstructions (Theis et al., 2022), even at ultra-low bit-rates. Unfortunately, the proposed method requires the communication of Gaussian samples across many steps, which remains intractable because the exponential runtime complexity of channel simulation (Goc & Flamich, 2024).
In this work, we take first steps towards a diffusion-based progressive codec that is computationally tractable. The key idea is to replace Gaussian distributions in the forward process with suitable uniform distributions and adjust the reverse process distributions accordingly. These modifications allow the application of universal quantization (Zamir & Feder, 1992) for simulating the uniform noise channel, avoiding the intractability of Gaussian channel simulation in (Theis et al., 2022).
Specifically, our contributions are as follows:
@@
    We investigate design choices of UQDM, specifying its forward and reverse processes largely by matching the moments of those in Gaussian diffusion, and obtain the best results when we learn the reverse-process variance as inspired by (Nichol & Dhariwal, 2021).
@@ 2 Background
Relative Entropy Coding (REC) deals with the problem of efficiently communicating a single sample from a target distribution $q$ using a coding distribution $p$. Suppose two parties in communication have access to a common “prior” distribution $p$ and pseudo-random number generators with a common seed; a Relative Entropy Coding (REC) method (Flamich et al., 2020) allows the sender to transmit a sample ${\mathbf{z}}\sim q$ using close to $\mathrm{KL}(q\,\|\,p)$ bits on average. If $q$ arises from a conditional distribution, e.g., $q_{\mathbf{x}}=q({\mathbf{z}}\mid{\mathbf{x}})$ is the inference distribution of a VAE (which can be viewed as a noisy channel), a reverse channel coding or channel simulation (Theis & Ahmed, 2022) algorithm then allows the sender to transmit ${\mathbf{z}}\sim q_{\mathbf{x}}$ with ${\mathbf{x}}\sim p({\mathbf{x}})$ using close to $\operatorname{\mathbb{E}}_{{\mathbf{x}}\sim p({\mathbf{x}})}[\mathrm{KL}(q({\mathbf{z}}\mid{\mathbf{x}})\,\|\,p({\mathbf{z}}))]$ bits on average. At a high level, a typical REC method works as follows. The sender generates a (possibly large) number of candidate ${\mathbf{z}}$ samples from the prior $p$,
@@
A major challenge of REC algorithms is that their computational complexity generally scales exponentially with the amount of information being communicated (Agustsson & Theis, 2020; Goc & Flamich, 2024). As an example, the MRC algorithm (Cuff, 2008; Havasi et al., 2018) draws $M$ candidate samples and selects $K\in\{1,2,,...,M\}$ with a probability proportional to the importance weights, $\nicefrac{{q({\mathbf{z}}_{n})}}{{p({\mathbf{z}}_{n})}},n=1,...,M$; similarly to importance sampling, $M$ needs to be on the order of $2^{\operatorname{KL}(q\|p)}$ for ${\mathbf{z}}_{K}$ to be (approximately) a fair sample from $q$, thus requiring a number of drawn samples that scales exponentially with the relative entropy $\operatorname{KL}(q\|p)$ (the cost of transmitting $K$ is thus $\log M\approx\operatorname{KL}(q\|p)$ bits). The exponential complexity prevents, e.g., naively communicating the entire latent tensor ${\mathbf{z}}$ in a Gaussian VAE for lossy compression, as the relative entropy $\mathrm{KL}(q({\mathbf{z}}|{\mathbf{x}})\,\|\,p({\mathbf{z}}))$ easily exceeds thousands of bits, even for a small image. This difficulty can be partly remedied by performing REC on sub-problems with lower dimensions (Flamich et al., 2020; 2022) for which computationally viable REC algorithms exist (Flamich et al., 2024; Flamich, 2024), but at the expense of worse bitrate efficiency due to the accumulation of codelength overhead across the dimensions.
@@
In the above, the optimal entropy coding distribution $\mathbb{P}(K|U^{\prime}=u^{\prime})$ is obtained by discretizing $p_{\tilde{Y}}:=p_{Y}\star\mathcal{U}(-\nicefrac{{\Delta}}{{2}},\nicefrac{{\Delta}}{{2}})$ on a grid of width $\Delta$ and offset by $U^{\prime}=u^{\prime}$ (Zamir & Feder, 1992), where $\star$ denotes convolution. If the true $p_{\tilde{Y}}$ is unknown, we can replace it with a surrogate density model $f_{\theta}(\tilde{y})$ during entropy coding and incur a higher coding cost,
@@ 5 Experiments
When considering the progressive compression performance of VDM and UQDM, we consider three ways of computing progressive reconstructions from ${\mathbf{z}}_{t}$: denoise, where $\hat{{\mathbf{x}}}=\hat{{\mathbf{x}}}_{\theta}({\mathbf{z}}_{t};t)$ is the prediction from the denoising network; ancestral, where $\hat{{\mathbf{x}}}\sim p({\mathbf{x}}|{\mathbf{z}}_{t})$ is drawn by ancestral sampling; and flow-based where $\hat{{\mathbf{x}}}\sim p({\mathbf{x}}|{\mathbf{z}}_{t})$ is computed deterministically using the probability flow ODE in (Theis et al., 2022). In Gaussian diffusion, the probability flow ODE produces the same trajectory of marginal distributions as ancestral sampling. In the case of UQDM, we apply the same update equations and observe similar benefits, likely due to the continuous-time equivalence of the underlying processes of UQDM and VDM. See Section B.3 for details. Note that DiffC-A and DiffC-F (Theis et al., 2022) directly correspond to our VDM results with ancestral and flow-based reconstructions.
In all experiments involving VDM and UQDM, we always use the same denoising U-net architecture for both, except UQDM uses twice as many output dimensions to additionally predict the reverse-process variance (see Sec. 3). We refer to Appendix Sec. C for further experiment details.
@@ 5.1 Swirl Toy Data
Finally, we examine the quality of unconditional samples from UQDM with varying $T$. Although our earlier results indicate worse compression performance for $T>5$, Figure 7 shows that UQDM’s sample quality monotonically improves with increasing $T$.
@@ 5.2 CIFAR10
Next, we apply our method to natural images. We start with the CIFAR10 dataset containing $32\times 32$ images. We train a baseline VDM model with a smaller architecture than that used by Kingma et al. (2021), converging to around 3 bits per dimension. We use the noise schedule $\sigma_{t}^{2}=\mathrm{\sigma}(\gamma_{t})$ where $\gamma_{t}$ is linear in $t$ with learned endpoints $\gamma_{T}$ and $\gamma_{0}$. For our UQDM model we empirically find that $T\approx 4$ yields the best trade-off between bit-rate and reconstruction quality. We train our model end-to-end on the progressive coding NELBO eq. (1) with learned reverse-process variances.
@@
As shown in Figure 3, we consistently outperform both JPEG and JPEG2000 over all bit-rates and metrics. Even though BPG, a competitive non-progressive codec optimized for rate-distortion performance, achieves better reconstruction fidelity (as measured in PSNR) in the low bit-rate regime, our method closely matches BPG in realism (as measured in FID) and even beats BPG in PSNR at higher bit-rates. The theoretical performance of compression with Gaussian diffusion (e.g., VDM) (Theis et al., 2022), especially with a high number of steps such as $T=1000$, is currently computationally infeasible, both due to the large number of neural function evaluations required, and due the intractable runtime of REC algorithms in the Gaussian case. Still, for reference we report theoretical results both for $T=1000$ and $T=20$, where the latter uses a smaller and more practical number of diffusion/progressive reconstruction steps.
@@ 6 Discussion
Future work may further improve our approach to close the performance gap to Gaussian diffusion; the latter represents the ideal lossy compression performance under a perfect realism constraint for an approximately Gaussian-distributed data source (Theis et al., 2022). This may require more sophisticated methods for computing progressive reconstructions that can achieve higher quality with fewer steps, or exploring different parameterizations of the forward and reverse process with better theoretical properties. Finally, we expect further improvement in computation efficiency and scalability when combining our method with ideas such as latent diffusion (Rombach et al., 2022), distillation (Sauer et al., 2024), or consistency models (Song et al., 2023).
@@ Acknowledgments
Justus Will and Yibo Yang acknowledge support from the HPI Research Center in Machine Learning and Data Science at UC Irvine. Stephan Mandt acknowledges support from the National Science Foundation (NSF) under an NSF CAREER Award IIS-2047418 and IIS-2007719, the NSF LEAP Center, by the Department of Energy under grant DE-SC0022331, the IARPA WRIVA program, the Hasso Plattner Research Center at UCI, the Chan Zuckerberg Initiative, and gifts from Qualcomm and Disney. We thank Kushagra Pandey for feedback on the manuscript.
@@ A.2 Uniform (Ours)
${\mathbf{z}}_{t}=b_{T|t}\,{\mathbf{z}}_{T}+c_{T|t}\,{\mathbf{x}}+\underbrace{\sum_{v=t+1}^{T}\sqrt{12}\delta_{v|t}\mathbf{u}_{v}}_{\coloneqq{\boldsymbol{\omega}}_{t}},$ (8)
@@
$\delta_{v|t}:=\beta_{v|v-1}\prod_{j=t+1}^{v-1}b_{j|j-1}=\frac{\sigma_{t}^{2}}{\alpha_{t}}\sqrt{\mathrm{SNR}(v-1)-\mathrm{SNR}(v)},$
where
$\mathrm{SNR}(s):=\frac{\alpha_{s}^{2}}{\sigma_{s}^{2}}.$
@@
$\mathrm{Var}\left(\boldsymbol{\omega}_{t}\right)$ $=\sum_{v=t+1}^{T}\delta_{v|t}^{2}\mathbf{I}=\frac{\sigma_{t}^{4}}{\alpha^{2}_{t}}[\mathrm{SNR}(t)-\mathrm{SNR}(T)]\mathbf{I}=\beta^{2}_{T|t}\mathbf{I},$
@@ A.3 Convergence to the Gaussian case
We show that both forward processes are equivalent in the continuous-time limit. To allow comparison across different number of steps $T$, we suppose that $\alpha_{t}$ and $\sigma_{t}$ are obtained from continuous-time schedules $\alpha(\cdot):[0,1]\to\mathbb{R}^{+}$ and $\sigma(\cdot):[0,1]\to\mathbb{R}^{+}$ (which were fixed ahead of time), such that $\alpha_{t}:=\alpha(t/T)$ and $\sigma_{t}:=\sigma(t/T)$ for $t=0,\dots,T$, for any choice of $T$. As in VDM (Kingma et al., 2021), we assume that the continuous-time signal-to-noise ratio $\mathrm{snr}(\cdot):=\alpha(\cdot)^{2}/\sigma(\cdot)^{2}$ is strictly monotonically decreasing.
To obtain the continuous-time limit, we hold the “continuous” time $\rho:=\frac{t}{T}$ fixed for some $\rho\in[0,1)$, and let $T\to\infty$ (or equivalently, let the time discretization $\frac{1}{T}\to 0$). We note that the quantities $b_{T|t}$, $c_{T|t}$, $\beta^{2}_{T|t}$ only depend on $\rho$, and are thus well-defined when we hold $\rho$ fixed and let $T\to\infty$:
$b_{T|t}$ $=\frac{\alpha_{T}}{\alpha_{t}}\frac{\sigma_{t}^{2}}{\sigma_{T}^{2}}=\frac{\alpha(1)}{\alpha(\rho)}\frac{\sigma^{2}(\rho)}{\sigma^{2}(1)},$
$c_{T|t}$ $=\left(\sigma^{2}(1)-\frac{\alpha^{2}(1)}{\alpha^{2}(\rho)}\sigma^{2}(\rho)\right)\frac{\alpha(\rho)}{\sigma^{2}(1)},$
$\beta^{2}_{T|t}$ $=\left(\sigma^{2}(1)-\frac{\alpha^{2}(1)}{\alpha^{2}(\rho)}\sigma^{2}(\rho)\right)\frac{\sigma^{2}(\rho)}{\sigma^{2}(1)}=\frac{\sigma^{4}(\rho)}{\alpha^{2}(\rho)}(\mathrm{snr}(\rho)-\mathrm{snr}(1)).$
We start by showing that our $q({\mathbf{z}}_{t}|{\mathbf{z}}_{T},{\mathbf{x}})$ converges to the corresponding Gaussian distribution in VDM in the continuous-time limit, which in turn implies the convergence of our $q({\mathbf{z}}_{t}|{\mathbf{x}})$ to the corresponding Gaussian distribution in VDM.
@@
Recall the following fact in the forward process of UQDM (see eq. (8)):
${\mathbf{z}}_{t}=b_{T|t}\,{\mathbf{z}}_{T}+c_{T|t}\,{\mathbf{x}}+\underbrace{\sum_{v=t+1}^{T}\sqrt{12}\delta_{v|t}\mathbf{u}_{v}}_{\coloneqq{\boldsymbol{\omega}}_{t}},$ (9)
where
are independent uniform noise variables, and
$\delta_{v|t}:=\beta_{v|v-1}\prod_{j=t+1}^{v-1}b_{j|j-1}=\frac{\sigma_{t}^{2}}{\alpha_{t}}\sqrt{\mathrm{SNR}(v-1)-\mathrm{SNR}(v)},$
where
$\mathrm{SNR}(s):=\frac{\alpha_{s}^{2}}{\sigma_{s}^{2}}.$
It therefore suffices to show that $\boldsymbol{\omega}_{t}$ converges in distribution to $\mathcal{N}(\mathbf{0},\beta^{2}_{T|t}\mathbf{I})$ in the continuous-time limit. Since the different coordinates of $\boldsymbol{\omega}_{t}$ are independent, we focus on a single coordinate and study the continuous-time limit of a scalar $\Omega_{t}$, given by a sum of scaled uniform variables,
$\Omega_{t}$ $:=\sum_{v=t+1}^{T}\left(\frac{\sqrt{12}\sigma^{2}(\rho)}{\alpha(\rho)}\sqrt{\mathrm{snr}(\frac{v-1}{T})-\mathrm{snr}(\frac{v}{T})}\right)U_{v}$ (10)
$=\sum_{j=1}^{n}\left(\frac{\sqrt{12}\sigma^{2}(\rho)}{\alpha(\rho)}\sqrt{\mathrm{snr}(\rho+\frac{j-1}{T})-\mathrm{snr}(\rho+\frac{j}{T})}\right)U_{j}$ (11)
where $U_{j}$’s are i.i.d. $\mathcal{U}(-\nicefrac{{1}}{{2}},\nicefrac{{1}}{{2}})$ variables, and in the last step we set $n:=n(T)=T-t$ and switched the summation index to $j=v-t$.
Define a triangular array of variables by
$X_{n,j}=\left(\frac{\sqrt{12}\sigma^{2}(\rho)}{\alpha(\rho)}\sqrt{\mathrm{snr}(\rho+\frac{j-1}{T})-\mathrm{snr}(\rho+\frac{j}{T})}\right)U_{j},$
for $j=1,2,...,n$ and for $n\in\mathbb{N}^{+}$. For each $n$, $\{X_{n,j}\}_{j=1,2,...,n}$ are independent variables with $\mathbb{E}[X_{n,j}]=0$, and it can be verified that
$\sum_{j=1}^{n}\mathbb{E}[X_{n,j}^{2}]=\mathrm{Var}\left(\Omega_{t}\right)=\beta^{2}_{T|t}=\frac{\sigma^{4}(\rho)}{\alpha^{2}(\rho)}(\mathrm{snr}(\rho)-\mathrm{snr}(1)).$
To apply the Lindeberg-Feller central limit theorem (Durrett, 2019, Theorem 3.4.10) to $\Omega_{t}=X_{n,1}+...+X_{n,n}$, it remains to verify the condition
$\forall\epsilon>0,\lim_{n\to\infty}\sum_{j=1}^{n}\mathbb{E}[X_{n,j}^{2}\mathbf{1}\{\left|X_{n,j}\right|>{\epsilon}\}]=0.$
Let $\epsilon>0$. Since $\mathrm{snr}(\cdot)$ is continuous on a compact domain $[0,1]$, it is also uniformly continuous; then there exists a $\delta$ such that
$|\mathrm{snr}(x_{1})-\mathrm{snr}(x_{2})|<\left(\frac{\epsilon\alpha(\rho)}{\sqrt{12}\sigma^{2}(\rho)}\right)^{2},\quad\forall x_{1},x_{2},|x_{1}-x_{2}|<\delta.$ (12)
Let $T$ (and thus $n=T-t$) become sufficiently large such that $\frac{1}{T}<\delta$. Then, for all such $T$ (and thus $n$) sufficiently large, and for all $j$, it holds that $\mathbf{1}\{\left|X_{n,j}\right|>{\epsilon}\}=0$ almost everywhere:
$\mathbb{P}(|X_{n,j}|>\epsilon)$ $=\mathbb{P}\left(\left(\frac{\sqrt{12}\sigma^{2}(\rho)}{\alpha(\rho)}\sqrt{\mathrm{snr}(\rho+\frac{j-1}{T})-\mathrm{snr}(\rho+\frac{j}{T})}\right)|U_{j}|>\epsilon\right)$ (13)
$=\mathbb{P}\left(|U_{j}|>\frac{\epsilon\alpha(\rho)}{\sqrt{12}\sigma^{2}(\rho)}\frac{1}{\sqrt{\mathrm{snr}(\rho+\frac{j-1}{T})-\mathrm{snr}(\rho+\frac{j}{T})}}\right)$ (14)
@@
$=0$ (16)
since $U_{j}\sim\mathcal{U}(-\nicefrac{{1}}{{2}},\nicefrac{{1}}{{2}})$, and it follows that
$\mathbb{E}[X_{n,j}^{2}\mathbf{1}\{\left|X_{n,j}\right|>{\epsilon}\}]=0$
for all $j$ for all sufficiently large $n$. We conclude by the Lindeberg-Feller theorem that
[…]
as $T\to\infty$. Applying the above argument coordinate-wise then proves the original statement. ∎
@@
As $q({\mathbf{z}}_{T}|x)=\mathcal{N}(\alpha_{T}{\mathbf{x}},\sigma^{2}_{T}\mathbf{I})$ does not depend on $T$, the joint distribution $q({\mathbf{z}}_{t},{\mathbf{z}}_{T}|x)=q({\mathbf{z}}_{t}|{\mathbf{z}}_{T},{\mathbf{x}})q({\mathbf{z}}_{T}|x)$ converges in distribution, which in turn implies convergence of $q({\mathbf{z}}_{t}|x)$. The statement then follows from the identity